# Sprecher Networks: A Parameter-Efficient Kolmogorov-Arnold Architecture

### v1

@@ 5 Parameter counting and efficiency as a trade-off
Sprecher Networks have a fundamentally different structure. Each block $\ell$ contains mixing weights $\lambda^{(\ell)}$ with $O(d_{\ell-1})$ parameters, where $d_{\ell-1}$ is the input dimension to that block (crucially, these are vectors, not matrices), shared splines $\phi^{(\ell)},\Phi^{(\ell)}$ with $2\times O(G)=O(G)$ parameters per block independent of $N$, and a shift parameter $\eta^{(\ell)}$ with $O(1)$ parameter per block. When lateral mixing is enabled, each block additionally contains a lateral scale parameter $\tau^{(\ell)}$ with $O(1)$ parameter and mixing weights $\omega^{(\ell)}$ with $O(d_{\ell})$ parameters (specifically $d_{\ell}$ for cyclic, $2d_{\ell}$ for bidirectional). When normalization is used, each block may have an associated normalization layer adding $O(N)$ parameters (specifically $2N$ for the affine transformation). When residual connections are used, cyclic residuals add at most $O(N)$ parameters per block (specifically $\max(d_{\ell-1},d_{\ell})$ for block $\ell$, or just 1 when dimensions match); a learned linear residual projection instead adds $O(d_{\ell-1}d_{\ell})$ parameters when $d_{\ell-1}\neq d_{\ell}$ (and just 1 when $d_{\ell-1}=d_{\ell}$). Summing over $L$ (or $L+1$) blocks, the total parameter count scales approximately as $O(LN+LG+L)$, where the linear dependence on $N$ now includes mixing weights, optional normalization parameters, lateral mixing weights, and residual connections.
@@
5.1 Illustrative parameter comparisons (hypothetical examples)
We present parameter count comparisons based on architectures reported in [13] to illustrate the potential parameter efficiency of SNs. Note that these examples are hypothetical; while the KAN architectures are from their paper, the SN results are theoretical projections that require empirical validation. Optimal architectures for SNs may differ from those of KANs due to the shared spline structure.
PDE solving example (KAN Ref § 3.4):
KAN architecture $[2,10,1]$, reported effective. $N_{in}=2,N_{hid}=10,N_{out}=1$.
  • KAN (est.): $(2\times 10+10\times 1)=30$ edges/splines. KANs use B-splines with $G_{\mathrm{KAN}}=20$ intervals, $k=3$ ($G_{\mathrm{KAN}}+k\approx 23$ params/spline). Total KAN params $\approx 30\times 23=690$.
  • SN (hypothetical): Architecture $2\to[10]\to 1$. $L=1$ hidden layer, scalar output means $L=1$ block total. Using piecewise linear splines with $G=21$ knots (20 intervals): $G=21$ params/spline. Shared splines: $2L\times G=2\times 21=42$. Mixing weights: $d_{0}=2$ (vector weights). Shift $\eta^{(1)}$: $1$. If lateral mixing enabled (cyclic): $\tau^{(1)}$: $1$, $\omega^{(1)}$: $10$. Total SN params $\approx 42+2+1+1+10=56$ (with lateral mixing) or $45$ (without).
  • Potential advantage: For equivalent structure, theoretical reduction factor $\approx 12-15$x. Actual performance requires empirical validation.
Knot theory example (KAN Ref § 4.3):
Pruned KAN $[17,1,14]$, with B-splines using $G_{\mathrm{KAN}}=3$ intervals, $k=3$ ($G_{\mathrm{KAN}}+k\approx 6$ params/spline).
  • KAN (est.): $(17\times 1+1\times 14)=31$ edges/splines. Total KAN params $\approx 31\times 6=186$.
  • SN (hypothetical): Architecture $17\to[1]\to 14$. $L=1$ hidden layer, vector output means $L+1=2$ blocks total. Using piecewise linear splines with $G=4$ knots (3 intervals): $2(L+1)\times G=4\times 4=16$. Mixing weights: $d_{0}+d_{1}=17+1=18$ (vector weights). Shifts $\eta^{(1)},\eta^{(2)}$: $2$. If lateral mixing enabled (cyclic) for both blocks: $\tau^{(1)},\tau^{(2)}$: $2$, $\omega^{(1)}$: $1$, $\omega^{(2)}$: $14$. Total SN params $\approx 16+18+2+2+1+14=53$ (with lateral mixing) or $36$ (without).
  • Potential advantage: Theoretical reduction factor $\approx 3.5-5$x. The narrow intermediate dimension $d_{1}=1$ might pose challenges for SN training.
These calculations illustrate the dramatic parameter reduction possible with SNs, but they also highlight a crucial practical consideration: the optimal architecture for SNs likely differs substantially from that of MLPs or KANs. With vector weights providing only linear scaling in width, SNs may require wider or deeper architectures to achieve comparable expressivity. The art of architecture selection for SNs involves balancing the parameter efficiency against the need for sufficient expressivity: a trade-off that remains poorly understood and likely depends strongly on the problem domain. Early empirical evidence suggests that SNs excel when the target function aligns well with their compositional, shift-based structure, but may require more iterations when forced to approximate functions that require truly independent processing of different output dimensions. The addition of lateral mixing partially addresses this limitation by enabling limited cross-dimensional communication.
@@ 6.1 Relation to Sprecher (1965) and universality
For any $n\geq 1$ and any $f\in C([0,1]^{n})$, there exists a single-layer SN with $\tau=0$ and no residual connection that represents $f$ exactly in the form (1) (taking $d_{\mathrm{out}}=2n+1$ and $\alpha=1$). In particular, for every $\epsilon>0$ there is a single-layer SN (with $\tau=0$ and no residual) such that $\sup_{\mathbf{x}\in[0,1]^{n}}|f(\mathbf{x})-\hat{f}(\mathbf{x})|<\epsilon$.
@@
Assume $\phi$ and $\Phi$ are Lipschitz on $I_{\phi}$ and $I_{\Phi}$ with constants $L_{\phi}:=\|\phi^{\prime}\|_{L^{\infty}(I_{\phi})}$ and $L_{\Phi}:=\|\Phi^{\prime}\|_{L^{\infty}(I_{\Phi})}$, respectively. Then $T$ is $L_{T}$-Lipschitz on $\mathcal{B}_{\mathrm{in}}$ in the $\ell_{\infty}$ norm with
@@
    $\phi^{(\ell)},\Phi^{(\ell)}\in C^{2}$ on neighborhoods of the intervals
@@ 7.1 Trainable splines
For practical implementations, we support both piecewise-linear splines and $C^{1}$ cubic Hermite splines with PCHIP (shape-preserving) slopes for both $\phi^{(\ell)}$ and $\Phi^{(\ell)}$. Piecewise-linear splines are the default when only function values are required (high efficiency, simple evaluation); the cubic option is used when higher-order derivatives must flow through the spline (e.g., the Poisson PINN benchmark, which requires second derivatives). In both cases, a spline with $G$ uniformly spaced knot locations (thus G1G-1 intervals) is parameterized by the values at these $G$ knot locations (thus $O(G)$ parameters per spline); for the cubic case, the PCHIP endpoint slopes are computed deterministically from these knot values. For range propagation, piecewise-linear extrema on an interval occur at the interval endpoints and/or knot points, whereas for cubic splines we additionally consider interior stationary points when computing tight bounds (see Lemma 4 and Algorithm 1). Each spline can be defined by a set of knots (x-coordinates) and corresponding coefficients (y-coordinates). A common approach is to use fixed, uniformly spaced knots over the spline’s domain, with learnable coefficients. The number of knots is a hyperparameter that affects both expressivity and computational cost.
For the inner spline $\phi^{(\ell)}$, we fix the codomain to $[0,1]$ and enforce strict monotonicity of the knot values by parameterizing positive differences. Concretely, let $(v_{k})_{k}$ be unconstrained parameters and set $\Delta_{k}=\operatorname{softplus}(v_{k})>0$; then the coefficients are
@@
with a tiny $\varepsilon>0$ (we use $\varepsilon=10^{-8}$) which yields strictly increasing knot values. Interpolating these knots with either piecewise–linear segments or a shape–preserving monotone $C^{1}$ cubic Hermite (PCHIP) scheme yields a monotone non-decreasing spline on its domain. Outside the domain we use constant extension, so $\phi^{(\ell)}(x)=0$ for inputs below the leftmost knot and $\phi^{(\ell)}(x)=1$ above the rightmost knot. The increments are initialized nearly uniform so that $\phi^{(\ell)}$ starts close to linear.
The outer spline $\Phi^{(\ell)}$ operates on a domain determined by the block’s bounds, and its codomain is parameterized as an interval $c_{c}^{(\ell)}-c_{r}^{(\ell)},\,c_{c}^{(\ell)}+c_{r}^{(\ell)}\,]$ with trainable center $c_{c}^{(\ell)}$ and scale (initialized positive) $c_{r}^{(\ell)}$. We initialize $(c_{c}^{(\ell)},c_{r}^{(\ell)})$ from the computed input domain so that $\Phi^{(\ell)}$ is near identity at initialization, and the parameters then adapt the output range during training. Monotonicity is not required for $\Phi^{(\ell)}$.
@@ 7.2 Shifts, weights, and optimization
Each block includes the learnable scalar shift $\eta^{(\ell)}$ and the learnable mixing weight vector $\lambda^{(\ell)}\in\mathbb{R}^{d_{\ell-1}}$. While Sprecher’s original construction requires the shift parameter to be positive ($\eta>0$), practical implementations can relax this constraint. As shown in Lemma 4, the theoretical domain computation handles both positive and negative values of $\eta^{(\ell)}$, allowing it to be trained as an unconstrained parameter initialized to a small positive value.
@@
All learnable parameters (spline coefficients, $\eta^{(\ell)}$, $\lambda^{(\ell)}$, lateral mixing parameters, and potentially range parameters for $\Phi^{(\ell)}$) are trained jointly using gradient-based optimization methods like Adam [9] or LBFGS. The loss function is typically Mean Squared Error (MSE) for regression tasks. Due to the constrained parameter space and shared spline structure, SNs may require more training iterations than equivalent MLPs or KANs to converge, though the per-iteration computational cost is typically lower due to fewer parameters.
@@ 7.3 Memory-efficient forward computation
A key advantage of SNs’ vector-based weight structure extends beyond parameter count to memory footprint, in particular to peak forward-intermediate memory. While dense MLP layers require $O(N^{2})$ parameters (and thus memory) for their weight matrices alone, SN blocks have only $O(N)$ vector parameters (plus $O(G)$ spline parameters) and can be evaluated without materializing the full $d_{\mathrm{in}}\times d_{\mathrm{out}}$ outer-product intermediate.
@@ 7.4 Theoretical domain computation
One advantage of the structured Sprecher architecture is the ability to compute provably safe interval bounds for spline inputs and outputs, which we use to adapt spline domains during training. For a fixed input interval to a spline, the exact range of a piecewise-linear spline on that interval is attained at the endpoints and any interior knots; for shape-preserving cubic (PCHIP) splines, stationary points within knot intervals must also be checked. When these per-spline range computations are propagated through multiple blocks using interval arithmetic, the resulting network-level bounds are generally conservative (since correlations between coordinates are ignored), but are typically much tighter than naive global bounds. Throughout this analysis, we assume inputs lie in $0,1]^{n}$, though the methodology extends naturally to any bounded domain. This assumption enables principled initialization and can aid in debugging and analysis.
@@
    A valid domain for $\Phi$ is therefore $\mathcal{D}_{\Phi}=\bigl[\min_{q}s_{q}^{\min,\mathrm{unmixed}},\;\max_{q}s_{q}^{\max,\mathrm{unmixed}}\bigr]$. (As a coarse bound, one may use only that $\phi(\cdot)\in[0,1]$ and thus replace the unknown values $\phi(a+\eta q)$ and $\phi(b+\eta q)$ by 0 and $1$, respectively.)
@@
    where $a^{+}=\max(a,0)$ and $a^{-}=\min(a,0)$. For bidirectional mixing, apply the analogous sign-split bound to each neighbor contribution $(\tau\omega_{q,\text{fwd}})\,s_{(q+1)\bmod d_{\mathrm{out}}}$ and $(\tau\omega_{q,\text{bwd}})\,s_{(q-1)\bmod d_{\mathrm{out}}}$, and sum the contributions.
@@
    Tight output bounds require checking all candidate extrema of $\Phi$ over the interval. For piecewise linear $\Phi$, extrema occur at endpoints and knot points, so it suffices to evaluate $\Phi$ at interval endpoints and all knots within the interval. For cubic $\Phi$ (PCHIP), extrema may also occur at internal critical points where $\Phi^{\prime}$ is zero, requiring evaluation at those points. (If an affine codomain reparameterization $(c_{c},c_{r})$ is used, it is an affine map of the underlying spline values; it does not change the set of candidate points for extrema, only rescales and shifts the resulting values.) This yields correct range propagation even for oscillatory $\Phi$.
@@
    • Broadcast: Each output $q$ receives $w_{q}^{\text{bcast}}\cdot x_{1+(q\bmod d_{\mathrm{in}})}$
@@
      where $t_{\min}=\min_{k}\tilde{\Phi}(u_{k})$ and $t_{\max}=\max_{k}\tilde{\Phi}(u_{k})$ are the minimum/maximum unnormalized knot values (with $\{u_{k}\}$ the knot locations) and $\epsilon>0$ is a small constant. The boundary derivatives are scaled by the same factor $2c_{r}/(t_{\max}-t_{\min}+\epsilon)$.
@@
This lemma enables a forward propagation algorithm for computing all spline domains throughout the network. The algorithm can optionally apply a small relative domain safety margin (e.g., 1%) to reduce edge hits during training; setting the margin to 0 yields the tightest domains. Crucially, the algorithm’s efficiency relies on the fact that spline extrema over compact intervals can be found by evaluating a small, explicit candidate set: for piecewise-linear splines, the interval endpoints and any knots inside the interval; for piecewise-cubic Hermite/PCHIP splines, the same candidates plus any interior critical points where the derivative vanishes (one quadratic solve per segment).
@@
Figure 7: Spline domain resampling during training (applied to the outer spline $\Phi$). As parameters $\lambda$ and $\eta$ evolve, induced spline input ranges shift dynamically. Left: Naive horizontal rescaling distorts the learned function by stretching and shifting the shape to fit the new domain. Right: Our resampling evaluates the previous spline at the new knot positions (for piecewise-linear $\Phi$, query points are clamped to the previous domain when a new knot lies outside it; for cubic/PCHIP $\Phi$, the previous spline’s linear extrapolation is used), approximately preserving the function values at the new knots that lie within the previous domain; the updated $\Phi$ still uses linear extrapolation outside its new domain.
@@
2:(Here $L_{\text{blocks}}$ is the number of Sprecher blocks: $L_{\text{blocks}}=L$ for scalar outputs, and $L_{\text{blocks}}=L{+}1$ for vector outputs.)
@@
7:   Apply normalization bounds to $\mathcal{R}_{\ell-1}$ (training: conservative shared bound derived from standardized [4,4][-4,4]; eval: use running stats and affine)
@@
17:  Evaluate $\Phi^{(\ell)}$ on each interval $s^{\min}_{q},s^{\max}_{q}]$ at (i) the endpoints, (ii) all $\Phi$-knots lying inside the interval, and (iii) for cubic/PCHIP $\Phi$, all interior critical points where $\frac{d}{dx}\Phi(x)=0$
@@
28:   Apply normalization bounds (training: conservative shared bound derived from standardized [4,4][-4,4]; eval: use running stats and affine)
@@
Remark 18 (Dynamic spline updates and function preservation).
A critical challenge in training Sprecher Networks is that the domains of the splines $\phi^{(\ell)}$ and $\Phi^{(\ell)}$ evolve as the parameters $\eta^{(\ell)}$, $\lambda^{(\ell)}$, and lateral mixing parameters are updated. To maintain theoretical fidelity to Sprecher’s formula while adapting to evolving domains, we employ a selective function-preserving resampling strategy:

### v2

@@ 5 Parameter counting and efficiency as a trade-off
Sprecher Networks have a fundamentally different structure. While a block can be viewed as densely connecting inputs to outputs computationally (as in Figure 2), SNs do not assign learnable parameters to individual edges: all input–output interactions share the same $\phi^{(\ell)}$ (and the same $\Phi^{(\ell)}$ after aggregation), with diversity introduced via the index shift and shared vector weights. Each block $\ell$ contains mixing weights $\lambda^{(\ell)}$ with $O(d_{\ell-1})$ parameters, where $d_{\ell-1}$ is the input dimension to that block (crucially, these are vectors, not matrices), shared splines $\phi^{(\ell)},\Phi^{(\ell)}$ with $2\times O(G)=O(G)$ parameters per block independent of $N$, and a shift parameter $\eta^{(\ell)}$ with $O(1)$ parameter per block. When lateral mixing is enabled, each block additionally contains a lateral scale parameter $\tau^{(\ell)}$ with $O(1)$ parameter and mixing weights $\omega^{(\ell)}$ with $O(d_{\ell})$ parameters (specifically $d_{\ell}$ for cyclic, $2d_{\ell}$ for bidirectional). When normalization is used, each block may have an associated normalization layer adding $O(N)$ parameters (specifically $2N$ for the affine transformation). When residual connections are used, cyclic residuals add at most $O(N)$ parameters per block (specifically $\max(d_{\ell-1},d_{\ell})$ for block $\ell$, or just 1 when dimensions match); a learned linear residual projection instead adds $O(d_{\ell-1}d_{\ell})$ parameters when $d_{\ell-1}\neq d_{\ell}$ (and just 1 when $d_{\ell-1}=d_{\ell}$). Summing over $L$ (or $L+1$) blocks, the total parameter count scales approximately as $O(LN+LG+L)$, where the linear dependence on $N$ now includes mixing weights, optional normalization parameters, lateral mixing weights, and residual connections.
@@ 6.1 Relation to Sprecher (1965) and universality
For any $n\geq 1$ and any $f\in C([0,1]^{n})$, there exists a single-layer SN with $\tau=0$ and no residual connection that represents $f$ exactly in the form (1) (taking $d_{\mathrm{out}}=2n+1$ and $\alpha=1$).
@@
Assume $\phi$ and $\Phi$ are Lipschitz on $I_{\phi}$ and $I_{\Phi}$ with Lipschitz constants $L_{\phi}$ and $L_{\Phi}$, respectively. Then $T$ is $L_{T}$-Lipschitz on $\mathcal{B}_{\mathrm{in}}$ in the $\ell_{\infty}$ norm with
@@
    The ideal constituent functions $\phi^{(\ell)}$ and $\Phi^{(\ell)}$ are in $C^{2}$ on neighborhoods of the intervals
@@
Remark 16 (Smoothness assumptions and the original Sprecher construction).
The approximation analysis above assumes that the ideal constituent functions $\phi^{(\ell)}$ and $\Phi^{(\ell)}$ are $C^{2}$ (and hence Lipschitz). This assumption is natural for our spline-parameterized networks, since piecewise-linear and PCHIP splines on bounded domains are automatically Lipschitz. However, it is worth noting that Sprecher’s original 1965 existence proof constructs an outer function $\Phi$ that is only guaranteed to be continuous, not necessarily Lipschitz or smooth. (The inner function $\phi$ in Sprecher’s construction is monotonic and Lipschitz by design.) Thus, while our approximation rates apply to the class of SNs parameterized by smooth splines, they do not directly characterize the regularity of the functions guaranteed to exist by Sprecher’s theorem. Extending the analysis to handle less regular $\Phi$ (e.g., merely continuous or Hölder continuous) remains an open theoretical question.
@@ 7.1 Trainable splines
For practical implementations, we support both piecewise-linear splines and $C^{1}$ cubic Hermite splines with PCHIP (Piecewise Cubic Hermite Interpolating Polynomial) slopes for both $\phi^{(\ell)}$ and $\Phi^{(\ell)}$. We use PCHIP (rather than, e.g., a B-spline basis) primarily for its shape-preservation: when the knot values are monotone it yields a monotone interpolant without oscillatory overshoot, which complements the monotonicity constraint used for $\phi^{(\ell)}$ in Sprecher Networks. Piecewise-linear splines are the default when only function values are required (high efficiency, simple evaluation); the cubic option is used when higher-order derivatives must flow through the spline—e.g., the Poisson PINN benchmark, which requires second derivatives (defined a.e. for $C^{1}$ Hermite splines). In both cases, a spline with $G$ uniformly spaced knot locations (thus G1G-1 intervals) is parameterized by the values at these $G$ knot locations (thus $O(G)$ parameters per spline); for the cubic case, the PCHIP knot slopes are computed deterministically from these knot values. This retains a simple “values-at-knots” parameterization, which we will also exploit for efficient range propagation (Lemma 4). For range propagation, piecewise-linear extrema on an interval occur at the interval endpoints and/or knot points. For cubic Hermite/PCHIP splines, extrema over an interval may occur at the interval endpoints, at interior knot points, or at interior critical points of the cubic segments (when present); we therefore evaluate all such candidates in the theoretical domain propagation of Section 7.5 to obtain sound and tight range bounds. The knot count $G$ is a hyperparameter that trades off expressivity and computational cost.
For the inner spline $\phi^{(\ell)}$, we fix the codomain to $[0,1]$ and enforce strict monotonicity of the knot values by parameterizing positive increments and normalizing. Concretely, let $(v_{k})_{k}$ be unconstrained parameters and set $\Delta_{k}=\operatorname{softplus}(v_{k})>0$; then define
@@
with a tiny $\varepsilon>0$ (we use $\varepsilon=10^{-8}$) to prevent division by zero; this yields strictly increasing knot values with $0<c_{0}<c_{1}<\cdots<c_{G-1}<1$ (and $c_{G-1}\to 1$ as $\varepsilon\to 0$). Interpolating these knots with either piecewise–linear segments or a shape–preserving monotone $C^{1}$ cubic Hermite (PCHIP) scheme yields a monotone non-decreasing spline on its domain. Many KAN implementations [12] parameterize edge activations using B-spline bases; we use piecewise-linear / PCHIP interpolation instead because $\phi$ is constrained to be monotone: with PCHIP, monotonicity can be enforced simply by requiring increasing knot values (as above), and the interpolation is shape-preserving (avoiding overshoot). Outside the domain we use constant extension, so $\phi^{(\ell)}(x)=0$ for inputs below the leftmost knot and $\phi^{(\ell)}(x)=1$ above the rightmost knot. The increments are initialized nearly uniform so that $\phi^{(\ell)}$ starts close to linear.
The outer spline $\Phi^{(\ell)}$ operates on a domain determined by the block’s bounds. Optionally, we parameterize its codomain via an interval $c_{c}^{(\ell)}-c_{r}^{(\ell)},\,c_{c}^{(\ell)}+c_{r}^{(\ell)}\,]$ with trainable center $c_{c}^{(\ell)}$ and scale (initialized positive) $c_{r}^{(\ell)}$; when this codomain parameterization is disabled, $\Phi^{(\ell)}$ is used directly without an explicit codomain normalization, so its output range is determined by its knot values. When enabled, we initialize $(c_{c}^{(\ell)},c_{r}^{(\ell)})$ from the computed $\Phi^{(\ell)}$ input domain (i.e., the pre-activation range) so that $\Phi^{(\ell)}$ is near identity at initialization, and the parameters then adapt the output range during training. Monotonicity is not required for $\Phi^{(\ell)}$.
@@ 7.2 Shifts, weights, and optimization
Each block includes the learnable scalar shift $\eta^{(\ell)}$ and the learnable mixing weight vector $\lambda^{(\ell)}\in\mathbb{R}^{d_{\ell-1}}$. A practical default is to initialize $\lambda^{(\ell)}$ with a variance-preserving scheme (e.g. i.i.d. $\mathcal{N}(0,2/d_{\ell-1})$) so that the pre-activations $s_{q}^{(\ell)}$ start at $O(1)$ scale, and to initialize $\eta^{(\ell)}$ to a small positive value on the order of $1/d_{\ell}$, so that the initial shifts $\eta^{(\ell)}q$ span an $O(1)$ range across $q=0,\dots,d_{\ell}-1$. While Sprecher’s original construction requires $\eta>0$, practical implementations can relax this constraint: the theoretical domain computation in Lemma 4 supports both positive and negative values of $\eta^{(\ell)}$, so we treat $\eta^{(\ell)}$ as an unconstrained parameter during training.
@@
All learnable parameters (spline coefficients, $\eta^{(\ell)}$, $\lambda^{(\ell)}$, lateral mixing parameters, and potentially range parameters for $\Phi^{(\ell)}$) are trained jointly using gradient-based optimization methods like Adam [9] or LBFGS. The loss function is typically Mean Squared Error (MSE) for regression tasks. Due to the constrained parameter space and shared spline structure, SNs may require more training iterations than equivalent MLPs or KANs to converge; per-iteration wall-clock cost is workload- and implementation-dependent (fewer parameters reduce optimizer overhead, but spline evaluation and memory-saving recomputation can increase runtime).
@@ 7.3 Memory-efficient forward computation
A key advantage of SNs’ vector-based weight structure extends beyond parameter count to memory footprint, in particular to peak forward-intermediate memory. This is particularly relevant when memory (not compute) is the primary constraint, e.g. on-device/edge inference, embedded accelerators, or training regimes where activation memory dominates. While dense MLP layers require $O(N^{2})$ parameters (and thus memory) for their weight matrices alone, SN blocks have only $O(N)$ vector parameters (plus $O(G)$ spline parameters) and can be evaluated without materializing the full $d_{\mathrm{in}}\times d_{\mathrm{out}}$ outer-product intermediate.
@@
7.4 Embedded device implementation
As a practical demonstration of SNs’ memory efficiency, we implement an SN for digit classification on a severely resource-constrained embedded device: a handheld gaming console with a 67 MHz ARM9 processor, no floating-point unit (FPU), and only 4 MB of RAM. Inference is performed via an emulator using Q16.16 fixed-point arithmetic.
[…]
We also train a deeper SN with architecture $784\to[1280,1280,1280]\to 10$, amounting to 16,594 parameters. An equivalent-depth MLP at width 1280 would require approximately 4.3 million parameters, exceeding the device’s available memory entirely. This demonstrates that the linear parameter scaling of SNs enables deployment of meaningfully expressive models on hardware where traditional dense architectures are simply infeasible.
Figure 8: Real-time inference on a resource-constrained embedded device (67 MHz ARM9 processor, no FPU, 4 MB RAM). An SN with 2,613 parameters classifies hand-drawn digits using Q16.16 fixed-point arithmetic. The top screen displays architecture information and softmax probabilities; the bottom touchscreen shows the user-drawn input digit.
@@ 7.5 Theoretical domain computation
One advantage of the structured Sprecher architecture is the ability to compute sound (typically conservative) interval bounds for spline inputs and outputs, which we use to adapt spline domains during training. When normalization layers are inserted between blocks, we propagate post-normalization bounds as described in Section 3.3 (for BatchNorm in batch-statistics mode these are heuristic, whereas in running-statistics mode the bounds are exact since the transform is affine). For a fixed input interval to a spline, the exact range of a piecewise-linear spline on that interval is attained at the endpoints and any interior knots; for cubic Hermite/PCHIP splines, we additionally evaluate any interior critical points of the cubic segments (when present) to obtain sound range bounds. When these per-spline range computations are propagated through multiple blocks using interval arithmetic, the resulting network-level bounds are generally conservative (since correlations between coordinates are ignored), but are typically much tighter than naive global bounds. Throughout this analysis, we assume inputs lie in $0,1]^{n}$, though the methodology extends naturally to any bounded domain. This assumption enables principled initialization and can aid in debugging and analysis.
@@
    A valid domain for $\Phi$ is therefore $\mathcal{D}_{\Phi}=\bigl[\min_{q}s_{q}^{\min,\mathrm{unmixed}},\;\max_{q}s_{q}^{\max,\mathrm{unmixed}}\bigr]$. (As a coarse bound, if $\phi$ is parameterized to satisfy $\phi(\cdot)\in[0,1]$, one may replace the unknown values $\phi(a+\eta q)$ and $\phi(b+\eta q)$ by 0 and $1$, respectively.)
@@
    where for any scalar $c$ we write $c^{+}=\max(c,0)$ and $c^{-}=\min(c,0)$. For bidirectional mixing, apply the analogous sign-split bound to each neighbor contribution $(\tau\omega_{q,\text{fwd}})\,s_{(q+1)\bmod d_{\mathrm{out}}}$ and $(\tau\omega_{q,\text{bwd}})\,s_{(q-1)\bmod d_{\mathrm{out}}}$, and sum the contributions.
@@
    Tight output bounds require checking all candidate extrema of $\Phi$ over the interval. For piecewise linear $\Phi$, extrema occur at endpoints and knot points, so it suffices to evaluate $\Phi$ at interval endpoints and all knots within the interval. For cubic $\Phi$ (PCHIP), the interpolant is shape-preserving (it does not overshoot the knot values). However, since a cubic segment can attain its extrema at interior critical points, tight range bounds require evaluating $\Phi$ at the interval endpoints, all knots within the interval, and any interior critical points that lie within the interval. (If an affine codomain reparameterization $(c_{c},c_{r})$ is used, it is an affine map of the underlying spline values; it does not change the set of candidate points for extrema, only rescales and shifts the resulting values.) This yields correct range propagation even for oscillatory $\Phi$.
@@
    • Broadcast: Each output $q$ receives $r_{q}=w_{q}^{\text{bcast}}\cdot x_{1+(q\bmod d_{\mathrm{in}})}$, hence
      $r_{q}\in\bigl[\min\{w_{q}^{\text{bcast}}a_{1+(q\bmod d_{\mathrm{in}})},w_{q}^{\text{bcast}}b_{1+(q\bmod d_{\mathrm{in}})}\},\max\{w_{q}^{\text{bcast}}a_{1+(q\bmod d_{\mathrm{in}})},w_{q}^{\text{bcast}}b_{1+(q\bmod d_{\mathrm{in}})}\}\bigr].$
@@
      where $t_{\min}=\min_{k}\tilde{\Phi}(x_{k})$ and $t_{\max}=\max_{k}\tilde{\Phi}(x_{k})$ are the minimum/maximum unnormalized knot values (with $\{x_{k}\}$ the knot locations) and $\epsilon>0$ is a small constant. The boundary derivatives are scaled by the same factor $2c_{r}/(t_{\max}-t_{\min}+\epsilon)$.
@@
This lemma enables a forward propagation algorithm for computing all spline domains throughout the network. The algorithm can optionally apply a small relative domain safety margin (e.g., 10%) to reduce edge hits during training; setting the margin to 0 yields the tightest domains. Crucially, the algorithm’s efficiency relies on the fact that spline extrema over compact intervals can be bounded by evaluating a small, explicit candidate set: for piecewise-linear splines, the interval endpoints and any knots inside the interval; for piecewise-cubic Hermite/PCHIP splines, the interval endpoints, knots inside the interval, and any interior critical points of the cubic pieces that lie inside the interval.
@@
Figure 9: Spline domain resampling during training (applied to the outer spline $\Phi$). As parameters $\lambda$ and $\eta$ evolve, induced spline input ranges shift dynamically. Left: Naive horizontal rescaling distorts the learned function by stretching and shifting the shape to fit the new domain. Right: Our resampling evaluates the previous spline at the new knot positions (for piecewise-linear $\Phi$, query points are clamped to the previous domain so any new knot outside the old domain uses the nearest boundary value; for cubic/PCHIP $\Phi$, out-of-domain query points are evaluated via the spline’s linear extrapolation determined by the endpoint derivative), approximately preserving the function values at the new knots that lie within the previous domain; the updated $\Phi$ still uses linear extrapolation outside its new domain.
@@
2:(Here $L_{\text{blocks}}$ is the number of Sprecher blocks: $L_{\text{blocks}}=L$ for the summed scalar-output form, and $L_{\text{blocks}}=L{+}1$ when an explicit non-summed output block is used (always for vector outputs, and optionally also for scalar outputs).)
@@
7:   Apply normalization bounds to $\mathcal{R}_{\ell-1}$ (batch-statistics: conservative shared bound derived from standardized [4,4][-4,4]; running-statistics: use running stats and affine)
@@
17:  Evaluate $\Phi^{(\ell)}$ on each interval $s^{\min}_{q},s^{\max}_{q}]$ at (i) the endpoints, (ii) all $\Phi$-knots lying inside the interval, and (iii) (for cubic $\Phi$) any interior critical points of the relevant cubic pieces lying inside the interval
@@
28:   Apply normalization bounds (batch-statistics: heuristic shared bound based on standardized [4,4][-4,4]; running-statistics: use running stats and affine)
@@
Remark 19 (Dynamic spline updates and value-preserving resampling).
A critical challenge in training Sprecher Networks is that the domains of the splines $\phi^{(\ell)}$ and $\Phi^{(\ell)}$ evolve as the parameters $\eta^{(\ell)}$, $\lambda^{(\ell)}$, and lateral mixing parameters are updated. To maintain theoretical fidelity to Sprecher’s formula while adapting to evolving domains, we employ a selective value-preserving resampling strategy: